# Understanding Structured Financial Data with LLMs: A Case Study on Fraud Detection

### v1

@@ 1 Introduction
Financial fraud is a pervasive and evolving threat that costs businesses and consumers billions of dollars every year hilal2022financial. The fraud landscape is highly dynamic, with malicious actors continually devising new strategies to circumvent existing safeguards. In response to the limitations of rule-based systems soui2019rule, machine learning (ML) models have been adopted to identify illicit or abnormal transactions from large-scale financial data jin2025analysis. However, their performance is critically dependent on extensive and laborious feature engineering. This reliance makes them costly to develop and maintain. Furthermore, these models are sensitive to pattern shifts gama2014survey, where the statistical properties of the data change over time, requiring frequent retraining and model updates. These trained models also offer limited interpretability beyond aggregate feature importance scores, posing challenges to transparent decision-making.
Recent advances in LLMs offer a promising alternative across many domains wang-etal-2025-user; hu2024enhancing; ning2025dima; wang2024biorag, including finance wu2023bloomberggpt; wang2023fingpt; yu2024fincon. With strong capabilities in complex reasoning and emulating the logic of human analysts brown2020language, LLMs hold the potential to not only detect fraud but also provide human-readable rationales for their decisions. However, prior studies typically use LLMs as auxiliary tools to augment traditional classifiers rather than as standalone detectors. For example, yang2025flag and huang2025can integrate LLMs to improve the performance of Graph Neural Networks in fraud detection. Nevertheless, applying LLMs directly to fraud detection remains challenging. Benchmarks feng2023empowering; xie2024finben highlight that existing LLMs still struggle with financial risk prediction, achieving only marginal improvements over random guessing.
@@
  • Tabular Input Misalignment: Unlike frauds such as internet scams or phone scams that involve rich text data yang2025fraud; singh2025advanced, financial transaction data are primarily numeric and categorical features in a tabular format. LLMs, being pretrained mainly on natural language, may not interpret the semantic meaning of structured features or handle numerical values with the required precision. In addition, real transactions can also involve hundreds or thousands of features, including all of them in a prompt may exceed context limits and introduce noise, as many are less significant signals. This lack of semantic context and high dimensionality makes it difficult for LLMs to detect fraud patterns from raw tabular data.
@@ 2.1 Fraud Detection
Traditional approaches for fraud detection relied heavily on rule-based systems and hand-crafted features, which, while interpretable, often struggled to adapt to evolving fraud patterns. In recent years, with the increasing availability of large-scale transaction data and the advances in deep learning, research has shifted towards more expressive models that capture more complex patterns. Traditional machine learning methods, particularly ensemble algorithms such as XGBoost chen2016xgboost, LightGBM ke2017lightgbm, and CatBoost prokhorenkova2018catboost, have shown strong effectiveness in capturing non-linear interactions among structured features. Beyond ensembles, deep neural networks have also been applied to detect different kinds of fraud fiore2019using; dou2020enhancing; li2024sefraud; yu2023group.
@@ 2.2 LLMs for finance
A growing line of research has emphasized the importance of benchmarking LLMs for financial applications. xie2023pixiu; xie2024finben build a large-scale benchmark covering a wide range of financial tasks, including fraud detection. InvestorBench li-etal-2025-investorbench evaluates LLM-based agents in financial decision-making settings, such as stock and cryptocurrency trading. Beyond text, FinMME luo-etal-2025-finmme and FCMR kim-etal-2025-fcmr extended the scope to multi-modal evaluation, incorporating tables, charts, and textual reports.
Parallel to benchmark construction, researchers have also introduced a series of domain-specific financial LLMs. Early efforts include FinBERT yang2020finbert, which adapted BERT for financial sentiment analysis; BloombergGPT wu2023bloomberggpt, a large-scale model trained on proprietary financial corpora; and FinGPT wang2023fingpt, which emphasizes open, low-cost, and continuously updated pipelines. Beyond general-purpose models, several frameworks have been designed for task-specific applications. For example, yu2024fincon proposed a multi-agent framework for stock trading and portfolio management, while aguda2024large explored leveraging LLMs as data annotators to extract relations in financial documents.
@@ 2.3 In-context learning and RAG
LLMs can adapt to new tasks at inference by conditioning on a few input–label examples in the prompt, without parameter updates brown2020language.This in-context learning (ICL) enables few-shot generalization through text alone. Prior work suggests that ICL may emulate implicit learning mechanisms, such as pattern matching or regression, while remaining sensitive to the choice and order of demonstrations min2022rethinking.
Retrieval-augmented generation (RAG) equips LLMs with external memory: a retriever selects relevant examples, and the generator conditions on them to produce grounded outputs lewis2020retrieval. The integration of external knowledge can enhance the accuracy of model responses and improve generation reliability guu2020retrieval; karpukhin2020dense; ram2023context. We extend this idea to structured data: retrieve semantically similar instances, present them as context, and let the model reason by comparison.
@@ 3.2 Offline Feature Reduction
Real-world transaction data often includes hundreds or even thousands of features, many of which carry less significant signal for fraud detection. If such high-dimensional data are fed directly into a retrieval-augmented LLM lewis2020retrieval, two problems arise: (i) long prompts exceed the model’s context window, making it infeasible to include all features and examples due to extra inference time and memory requirements, and (ii) irrelevant or noisy features dilute the predictive signal and hinder the model’s ability to learn meaningful patterns. Thus, in FinFRE-RAG, we consider reducing the feature space to ensure that the LLM focuses on the most informative attributes and that retrieval remains efficient.
@@ 3.3 Online Retrieval-Augmented In-Context Learning
The top-$n$ similar transactions constitute the retrieval set $\mathcal{D}_{\mathrm{retrieved}}(x_{q})$. The hybrid retrieval strategy combines semantic consistency (via categorical filtering) with graded similarity (via vector search), aligning the selection with the LLM’s strength at analogical reasoning yasunaga2024large.
@@ 4.1 Experiment Setup
Datasets. We evaluate on four public fraud detection datasets: ccf feng2023empowering, ccFraud feng2023empowering; kamaruddin2016credit, IEEE-CIS howard2019ieee, and PaySim lopez2017synthetic, covering real-world and synthetic transaction data with varying sizes, feature spaces, and fraud ratios. Detailed dataset descriptions are provided in the Appendix A. To ensure computational feasibility for inference while preserving class imbalance, we randomly sample 8,000 transactions for testing and 2,000 for validation, maintaining the original fraud ratio. The remaining transactions constitute the retrieval pool (external dataset $\mathcal{D}_{\mathrm{ext}}$) during inference.
Models. Considering the sensitive nature of financial data, which often must remain strictly in-house for compliance, we focus on open-weight instruction-tuned LLMs that can be deployed on-premise. Concretely, we evaluate Qwen3-14B and Qwen3-Next-80B-A3B-Thinking yang2025qwen3, Gemma 3-12B and Gemma 3-27B team2025gemma, as well as GPT-OSS-20B and GPT-OSS-120B agarwal2025gpt. These models represent a range of medium to large instruction-tuned LLMs. As non-LLM baselines for tabular fraud detection, we include Random Forest, XGBoost, and the recent deep tabular model TabM gorishniy2024tabm, covering widely used tree ensembles and a strong state-of-the-art neural baseline. There are also some domain‑specific financial LLMs, such as FinGPT yang2023fingpt and FinMA xie2023pixiu. However, these models are based on earlier Llama versions touvron2023llama2; touvron2023llama and are comparatively smaller, which may weaker instruction-following and reasoning for our setting, so we do not include them in experiments.
Evaluation Metrics. Due to the class imbalance, accuracy is not informative. We therefore use F1-score and Matthews Correlation Coefficient (MCC) chicco2020advantages as primary metrics, as they better reflect performance under imbalance. We also report precision (share of flagged transactions that are truly fraudulent) and recall (share of frauds correctly identified) for analyst reference.
@@
We finetune Qwen3-14B, Gemma 3-12B, and GPT-OSS-20B via LoRA hu2022lora using unsloth unsloth, on the external split only. Since they are tuned on the external split, we do not apply RAG on these models; instead, we directly run inference for each transaction. We conduct a small grid over key hyperparameters on $10\%$ of the data to select rank, $\alpha$, learning rate, and learning rate schedulers, then train for one epoch. Full finetuning details (adapter targets, rank, learning rates, and batch size) are in Appendix B.
@@ 4.2 Results
Baseline performance. The results for baseline LLMs are consistent with findings in xie2024finben. LLMs perform poorly when directly prompted to classify fraud data without additional context, and they are substantially below strong tabular methods like TabM. For instance, Qwen3‑14B achieves a negative MCC score on the ccf, indicating that the model’s predictions on this dataset are indistinguishable from random guessing. Similar trends hold for other models. These results confirm that LLMs, when applied to challenging financial tabular data, fail to capture fraud patterns.
@@ Appendix A Datasets
Dataset
# Numerical
Features
# Categorical
Features
# Positive
# Negative
Imbalanced
License
@@
  • ccf [^1] https://www.kaggle.com/datasets/mlg-ulb/creditcardfraud feng2023empowering: A credit card fraud dataset with 284,807 transactions from European cardholders in September 2013. Each record includes 30 features, 28 of which are anonymized via PCA for confidentiality. The dataset is highly imbalanced, with only 0.172% fraudulent cases.
  • ccFraud [^2] https://github.com/The-FinAI/CALM/blob/main/data/fraud detection/ccFraud/ feng2023empowering; kamaruddin2016credit: It is a simulated dataset contains about 1 million transactions with 7 features (e.g., gender, account balance, transaction count). The proportion of fraudulent samples is 5.98%.
  • IEEE-CIS [^3] https://www.kaggle.com/competitions/ieee-fraud-detection/overview howard2019ieee: IEEE-CIS is a real-world dataset. We use the transaction-level data, which consists of 590,540 transactions with hundreds of anonymized numerical and categorical features. Fraudulent transactions make up 3.5% of the data, presenting challenges in both dimensionality and class imbalance. The license can be found on Kaggle  [^4] https://www.kaggle.com/competitions/ieee-fraud-detection/rules.
  • PaySim [^5] https://www.kaggle.com/datasets/ealaxi/paysim1 lopez2017synthetic: A large-scale simulated mobile money transaction dataset with 6.3 million samples. Each transaction is labeled as fraud or non-fraud, with features such as transaction type, amount, and balance updates. The fraud ratio is 0.13%. As noted in the dataset documentation, some columns (oldbalanceOrg, newbalanceOrig, oldbalanceDest, newbalanceDest) can lead to information leakage, as models may exploit them as shortcuts for fraud detection. We nevertheless include these features, since removing them would leave very limited information for learning. Therefore, evaluation results on PaySim should be interpreted with caution, and more emphasis should be placed on the other three datasets.
@@ Appendix B Hyperparameters
We tune all supervised baselines with Optuna akiba2019optuna, using 50 trials per dataset–model pair. Each trial is trained on $\mathcal{D}_{\mathrm{ext}}$; the configuration with the best validation MCC is selected and evaluated on the test set. For TabM training, we deployed early stopping with patience of 5 epochs. For tree methods, we set class weights via scale_pos_weight (XGBoost) or class_weight=balanced (RF) to mitigate the impact of class imbalance. Table 6 lists the search spaces.
@@
We apply parameter-efficient finetuning (PEFT) with LoRA hu2022lora to Qwen3-14B, Gemma 3-12B, and GPT-OSS-20B using Unsloth unsloth. To control computation cost, we conduct a small grid search on a 10% subsample of the external split ($\mathcal{D}_{ext}$) and select the configuration with the best validation loss. The selected configuration is then trained on $D_{ext}$ for one epoch. Our grid varies only the learning rate, scheduler, and LoRA rank; all other settings are fixed for stability and comparability. We always use 10% iterations for the warm-up.

### v2

@@ 1 Introduction
Financial fraud is a pervasive and evolving threat that costs businesses and consumers billions of dollars every year Hilal et al. (2022). The fraud landscape is highly dynamic, with malicious actors continually devising new strategies to circumvent existing safeguards. In response to the limitations of rule-based systems Soui et al. (2019), machine learning (ML) models have been adopted to identify illicit or abnormal transactions from large-scale financial data Jin and Zhang (2025). However, their performance is critically dependent on extensive and laborious feature engineering. This reliance makes them costly to develop and maintain. Furthermore, these models are sensitive to pattern shifts Gama et al. (2014), where the statistical properties of the data change over time, requiring frequent retraining and model updates. These trained models also offer limited interpretability beyond aggregate feature importance scores, posing challenges to transparent decision-making.
Recent advances in LLMs offer a promising alternative across many domains Wang et al. (2025); Hu et al. (2024); Ning et al. (2025); Wang et al. (2024); Tan et al. (2026), including finance Wu et al. (2023); Wang et al. ; Yu et al. (2024). With strong capabilities in complex reasoning and emulating the logic of human analysts Brown et al. (2020), LLMs hold the potential to not only detect fraud but also provide human-readable rationales for their decisions. However, prior studies typically use LLMs as auxiliary tools to augment traditional classifiers rather than as standalone detectors. For example, Yang et al. (2025b) and Huang and Wang (2025) integrate LLMs to improve the performance of Graph Neural Networks in fraud detection. Nevertheless, applying LLMs directly to fraud detection remains challenging. Benchmarks Feng et al. (2023); Xie et al. (2024) highlight that existing LLMs still struggle with financial risk prediction, achieving only marginal improvements over random guessing.
@@
  • Tabular Input Misalignment: Unlike frauds such as internet scams or phone scams that involve rich text data Yang et al. (2025c); Singh et al. (2025), financial transaction data are primarily numeric and categorical features in a tabular format. LLMs, being pretrained mainly on natural language, may not interpret the semantic meaning of structured features or handle numerical values with the required precision. In addition, real transactions can also involve hundreds or thousands of features, including all of them in a prompt may exceed context limits and introduce noise, as many are less significant signals. This lack of semantic context and high dimensionality makes it difficult for LLMs to detect fraud patterns from raw tabular data.
@@ 2.1 Fraud Detection
Traditional approaches for fraud detection relied heavily on rule-based systems and hand-crafted features, which, while interpretable, often struggled to adapt to evolving fraud patterns. In recent years, with the increasing availability of large-scale transaction data and the advances in deep learning, research has shifted towards more expressive models that capture more complex patterns. Traditional machine learning methods, particularly ensemble algorithms such as XGBoost Chen and Guestrin (2016), LightGBM Ke et al. (2017), and CatBoost Prokhorenkova et al. (2018), have shown strong effectiveness in capturing non-linear interactions among structured features. Beyond ensembles, deep neural networks have also been applied to detect different kinds of fraud Fiore et al. (2019); Dou et al. (2020); Li et al. (2024); Yu et al. (2023).
@@ 2.2 LLMs for finance
A growing line of research has emphasized the importance of benchmarking LLMs for financial applications. Xie et al. (2023, 2024) build a large-scale benchmark covering a wide range of financial tasks, including fraud detection. InvestorBench Li et al. (2025) evaluates LLM-based agents in financial decision-making settings, such as stock and cryptocurrency trading. Beyond text, FinMME Luo et al. (2025) and FCMR Kim et al. (2025) extended the scope to multi-modal evaluation, incorporating tables, charts, and textual reports.
Parallel to benchmark construction, researchers have also introduced a series of domain-specific financial LLMs. Early efforts include FinBERT Yang et al. (2020), which adapted BERT for financial sentiment analysis; BloombergGPT Wu et al. (2023), a large-scale model trained on proprietary financial corpora; and FinGPT Wang et al. , which emphasizes open, low-cost, and continuously updated pipelines. Beyond general-purpose models, several frameworks have been designed for task-specific applications. For example, Yu et al. (2024) proposed a multi-agent framework for stock trading and portfolio management, while Aguda et al. (2024) explored leveraging LLMs as data annotators to extract relations in financial documents.
@@ 2.3 In-context learning and RAG
LLMs can adapt to new tasks at inference by conditioning on a few input–label examples in the prompt, without parameter updates Brown et al. (2020).This in-context learning (ICL) enables few-shot generalization through text alone. Prior work suggests that ICL may emulate implicit learning mechanisms, such as pattern matching or regression, while remaining sensitive to the choice and order of demonstrations Min et al. (2022).
Retrieval-augmented generation (RAG) equips LLMs with external memory: a retriever selects relevant examples, and the generator conditions on them to produce grounded outputs Lewis et al. (2020). The integration of external knowledge can enhance the accuracy of model responses and improve generation reliability Guu et al. (2020); Karpukhin et al. (2020); Ram et al. (2023). We extend this idea to structured data: retrieve semantically similar instances, present them as context, and let the model reason by comparison.
@@ 3.2 Offline Feature Reduction
Real-world transaction data often includes hundreds or even thousands of features, many of which carry less significant signal for fraud detection. If such high-dimensional data are fed directly into a retrieval-augmented LLM Lewis et al. (2020), two problems arise: (i) long prompts exceed the model’s context window, making it infeasible to include all features and examples due to extra inference time and memory requirements, and (ii) irrelevant or noisy features dilute the predictive signal and hinder the model’s ability to learn meaningful patterns. Thus, in FinFRE-RAG, we consider reducing the feature space to ensure that the LLM focuses on the most informative attributes and that retrieval remains efficient.
@@ 3.3 Online Retrieval-Augmented In-Context Learning
The top-$n$ similar transactions constitute the retrieval set $\mathcal{D}_{\mathrm{retrieved}}(x_{q})$. The hybrid retrieval strategy combines semantic consistency (via categorical filtering) with graded similarity (via vector search), aligning the selection with the LLM’s strength at analogical reasoning Yasunaga et al. (2024).
@@ 4.1 Experiment Setup
Datasets. We evaluate on four public fraud detection datasets: ccf Feng et al. (2023), ccFraud Feng et al. (2023); Kamaruddin and Ravi (2016), IEEE-CIS Howard and Bouchon-Meunier (2019), and PaySim Lopez-Rojas (2017), covering real-world and synthetic transaction data with varying sizes, feature spaces, and fraud ratios. Detailed dataset descriptions are provided in the Appendix A. To ensure computational feasibility for inference while preserving class imbalance, we randomly sample 8,000 transactions for testing and 2,000 for validation, maintaining the original fraud ratio. The remaining transactions constitute the retrieval pool (external dataset $\mathcal{D}_{\mathrm{ext}}$) during inference.
Models. Considering the sensitive nature of financial data, which often must remain strictly in-house for compliance, we focus on open-weight instruction-tuned LLMs that can be deployed on-premise. Concretely, we evaluate Qwen3-14B and Qwen3-Next-80B-A3B-Thinking Yang et al. (2025a), Gemma 3-12B and Gemma 3-27B Team et al. (2025), as well as GPT-OSS-20B and GPT-OSS-120B Agarwal et al. (2025). These models represent a range of medium to large instruction-tuned LLMs. As non-LLM baselines for tabular fraud detection, we include Random Forest, XGBoost, and the recent deep tabular model TabM Gorishniy et al. (2024), covering widely used tree ensembles and a strong state-of-the-art neural baseline. There are also some domain‑specific financial LLMs, such as FinGPT Yang et al. (2023) and FinMA Xie et al. (2023). However, these models are based on earlier Llama versions Touvron et al. (2023b, a) and are comparatively smaller, which may weaker instruction-following and reasoning for our setting, so we do not include them in experiments.
Evaluation Metrics. Due to the class imbalance, accuracy is not informative. We therefore use F1-score and Matthews Correlation Coefficient (MCC) Chicco and Jurman (2020) as primary metrics, as they better reflect performance under imbalance. We also report precision (share of flagged transactions that are truly fraudulent) and recall (share of frauds correctly identified) for analyst reference.
@@
We finetune Qwen3-14B, Gemma 3-12B, and GPT-OSS-20B via LoRA Hu et al. (2022) using unsloth Daniel Han and team (2023), on the external split only. Since they are tuned on the external split, we do not apply RAG on these models; instead, we directly run inference for each transaction. We conduct a small grid over key hyperparameters on $10\%$ of the data to select rank, $\alpha$, learning rate, and learning rate schedulers, then train for one epoch. Full finetuning details (adapter targets, rank, learning rates, and batch size) are in Appendix B.
@@ 4.2 Results
Baseline performance. The results for baseline LLMs are consistent with findings in Xie et al. (2024). LLMs perform poorly when directly prompted to classify fraud data without additional context, and they are substantially below strong tabular methods like TabM. For instance, Qwen3‑14B achieves a negative MCC score on the ccf, indicating that the model’s predictions on this dataset are indistinguishable from random guessing. Similar trends hold for other models. These results confirm that LLMs, when applied to challenging financial tabular data, fail to capture fraud patterns.
@@ 5 Conclusion
Acknowledgements
This work was funded in part by the National Science Foundation under award numbers IIS-2202699, IIS-2416895, IIS-2301599, CMMI-2301601, and DMS-2529302.
@@ Appendix A Datasets
Dataset # Numerical Features # Categorical Features # Positive # Negative Imbalanced (%) License
@@
  • ccf [^1] https://www.kaggle.com/datasets/mlg-ulb/creditcardfraud Feng et al. (2023): A credit card fraud dataset with 284,807 transactions from European cardholders in September 2013. Each record includes 30 features, 28 of which are anonymized via PCA for confidentiality. The dataset is highly imbalanced, with only 0.172% fraudulent cases.
  • ccFraud [^2] https://github.com/The-FinAI/CALM/blob/main/data/fraud detection/ccFraud/ Feng et al. (2023); Kamaruddin and Ravi (2016): It is a simulated dataset contains about 1 million transactions with 7 features (e.g., gender, account balance, transaction count). The proportion of fraudulent samples is 5.98%.
  • IEEE-CIS [^3] https://www.kaggle.com/competitions/ieee-fraud-detection/overview Howard and Bouchon-Meunier (2019): IEEE-CIS is a real-world dataset. We use the transaction-level data, which consists of 590,540 transactions with hundreds of anonymized numerical and categorical features. Fraudulent transactions make up 3.5% of the data, presenting challenges in both dimensionality and class imbalance. The license can be found on Kaggle  [^4] https://www.kaggle.com/competitions/ieee-fraud-detection/rules.
  • PaySim [^5] https://www.kaggle.com/datasets/ealaxi/paysim1 Lopez-Rojas (2017): A large-scale simulated mobile money transaction dataset with 6.3 million samples. Each transaction is labeled as fraud or non-fraud, with features such as transaction type, amount, and balance updates. The fraud ratio is 0.13%. As noted in the dataset documentation, some columns (oldbalanceOrg, newbalanceOrig, oldbalanceDest, newbalanceDest) can lead to information leakage, as models may exploit them as shortcuts for fraud detection. We nevertheless include these features, since removing them would leave very limited information for learning. Therefore, evaluation results on PaySim should be interpreted with caution, and more emphasis should be placed on the other three datasets.
@@ Appendix B Hyperparameters
We tune all supervised baselines with Optuna Akiba et al. (2019), using 50 trials per dataset–model pair. Each trial is trained on $\mathcal{D}_{\mathrm{ext}}$; the configuration with the best validation MCC is selected and evaluated on the test set. For TabM training, we deployed early stopping with patience of 5 epochs. For tree methods, we set class weights via scale_pos_weight (XGBoost) or class_weight=balanced (RF) to mitigate the impact of class imbalance. Table 6 lists the search spaces.
@@
We apply parameter-efficient finetuning (PEFT) with LoRA Hu et al. (2022) to Qwen3-14B, Gemma 3-12B, and GPT-OSS-20B using Unsloth Daniel Han and team (2023). To control computation cost, we conduct a small grid search on a 10% subsample of the external split ($\mathcal{D}_{ext}$) and select the configuration with the best validation loss. The selected configuration is then trained on $D_{ext}$ for one epoch. Our grid varies only the learning rate, scheduler, and LoRA rank; all other settings are fixed for stability and comparability. We always use 10% iterations for the warm-up.